OpenStudy (lukecrayonz):

x=e^(3y), I just need the first step done. x=e^(3y), I just need the first step done. @Mathematics

jimthompson5910 (jim_thompson5910):

x=e^(3y) ln(x)=ln(e^(3y))

OpenStudy (lukecrayonz):

Uhhhh, and the next step please :S Is it ln(x)=3yln(e)? or eln(3y)

jimthompson5910 (jim_thompson5910):

ln(x)=3yln(e)

jimthompson5910 (jim_thompson5910):

you're pulling down the exponent and placing it as the outer coefficient